Given the demand function P=150 -0.5Q and the total Cost function TC=8Q

A. Maximize profit for a monopoly firm

B. Maximize profit for a perfectly competitive firm

C. Calculate the profit that monopolist lost to act as a perfectly competitive firm


1
Expert's answer
2021-12-19T18:10:38-0500

A. Maximize profit for a monopoly firm


The profit for the monopoly will be at the point where MR=MC


Given the above demand curve, and we know that the marginal revenue is twice as the demand curve, the marginal revenue is equal to

"MR=150-Q"


The marginal cost from the cost function is

"MC=8"

Equating MC and MR and solving for Q, we get

"150-Q=8\\[0.3cm]
Q=142"

The price that the monopoly will charge is equal to

"P=150-0.4(142)\\[0.3cm]
P=\$79"

The profit for the firm is equal to

"\rm Profit=(76-8)142=\$10,082"


B. Maximize profit for a perfectly competitive firm

For a perfectly competitive firm, profit will be maximized when P=MC. Therefore

"150-0.5Q=8\\[0.3cm]
0.5Q=142\\[0.3cm]
Q=284"

The maximum profit for perfectly competitive firm is equal to


"\rm Profit=(8-8)84=\$0"


C. Calculate the profit that monopolist lost to act as a perfectly competitive firm

The profit lost by the monopolist is equal to

"0-10082=\$10,082"